Solve the following equation. Be sure to check each answer in the original equation if you multiply both sides by an expression that contains the variable.


(x-5)/2 +(4x)/3 =7/6

Answer by RAY100(1637)

(x-5)/2 +(4x)/3 =7/6
.
{3(x-5) + 2(4x) } /6 = 7/6
.
3x-15 +8x =7
.
11x = 22
.
x=2
.
check,,-3/2 +8/3 =7/6
-1.5 +2.67 = 1.167
1.167 = 1.167,,,,,ok